What is the standard form of 1/3y=3/5x-4
and What is the standard form of y= 5/6x + 1/3
Thanks my name is Shanaynay and you are off the chain!

Answer by stanbon(75887)

What is the standard form of 1/3y=3/5x-4
Cross-multiply to get:
5x-4=9y
Put in the form ax+by=c
5x-9y=4
----------------
and What is the standard form of y= 5/6x + 1/3
Multiply thru by 6 to get:
6y=5x+2
Put in the form ax+by=c
5x-6y = -2
--------------
Cheers,
Stan H.